Solve the differential equation by variation of parameters.
4y'' − 8y' + 8y = e^x sec x

y(x) =______
1
Expert's answer
2020-07-01T17:13:25-0400

The given equation can be written as:

(4D2-8D+8)y = exsecx


complementary equation is:

4D2-8D+8=0

D=4+4i,4-4i

Since they are complex roots, general solution is

y= e4x(acos4x + bsin4x)

Particular solution by variation of parameters:

Let z=e4xcos4x and u= e4xsin4x

Wronskian(z,u) ="\begin{vmatrix}
 e^{4x}cos4x & e^{4x}sin4x \\
 4e^{4x}cos4x - 4e^{4x}sin4x & 4e^{4x}sin4x + 4e^{4x}cos4x
\end{vmatrix}"

=4e4x(sin24x+cos24x)=4e4x


Particular solution is

=-z("\int (u/W) dt+u(\int(z/W)dt"


= -e4xcos4x ("\int (e^{4x}sin4x/4e^{4x})" "+e^{4x}sin4x (\int(e^{4x}cos4x /4e^{4x})"

="-e^{4x}cos4x (-cos4x\/16) + e^{4x}sin4x (sin4x\/16)+c"

So the solution is

y=

e4x(acos4x + bsin4x)+ "-e^{4x}cos4x (-cos4x\/16) + e^{4x}sin4x (sin4x\/16)+c"